Hi! I'm supposed to write an equation for the line. the problem is Through (7,-9), perpendicular to 2x+7y=-49. Please help me!
Answer by mananth(16946)

2 x + 7 y = -49
Find the slope of this line

7 y = -2 x -49
Divide by 7
y = - 2/ 7 x -7
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 2/7

The slope of a line perpendicular to the above line will be the negative reciprocal 7/2
Because m1*m2 =-1
The slope of the required line will be 7/2

m= 7/2 ,point ( 7 , -9 )
Find b by plugging the values of m & the point in
y=mx+b
-9 = 49/ 2 + b
b= -67/2
m= 7/2
The required equation isy = (7/2)x-67/ 2
m.ananth@hotmail.ca